# The first phase of mass transfer in low-mass binaries: neither stable nor a common envelope

### v1

@@ Abstract
Context. The masses of the white dwarfs in a binary carry information about previous mass-transfer phases. The core mass – radius relation of low-mass giants gives the size of the progenitor of a helium white dwarf at the moment it last filled its Roche lobe. Previously, we used this information for a few observed systems to propose a new mass-transfer type, based on an angular momentum balance.
Aims. Our aim is to investigate if stable mass transfer instead of the angular momentum prescription is consistent with the observed double helium white dwarf masses.
Methods. We reconstruct the progenitor evolution of observed double helium white dwarfs using the core mass – radius relation and evaluate if the periods at the start of the second phases of mass transfer are consistent with the outcome of stable mass transfer. More generally, we calculate the mass distribution of double helium white dwarfs for three different progenitors scenarios: double common envelope (with parameter ), angular momentum prescription (with parameter ) and stable mass transfer.
Results. We find that the observed systems are generally not consistent with stable mass transfer. Stable mass transfer leads to a tight correlation between the two white dwarf masses in a binary that is not consistent with the observed mass distribution. Double common envelope evolution is a particularly poor fit to the observations. The angular momentum prescription can populate the observed mass distribution, but not perfectly.
Conclusions. We conclude that the first phase of mass transfer initiated on the red giant branch in low-mass systems does not generally proceed as stable mass transfer nor as common envelope, and thus is poorly understood. This may be related to the fact that for many observed binaries that have finished the first phase of mass transfer the orbit is eccentric, which is an unexpected outcome of mass transfer.
@@
Most stars in the universe have low masses and all of these low-mass stars (0.8-2) follow a similar evolution (e.g. Refsdal & Weigert 1971; Paczyński 1971): after the main sequence the helium core contracts, becomes degenerate and grows in mass via hydrogen shell burning. At the same time, the envelope of the star expands, and the star ascends the giant branch, until the core experiences the helium flash around a mass of 0.47. The expansion of these stars happens for a significant part when they have deep convective envelopes.
@@
For a long time, it was believed that the first mass transfer phase for binary systems with such stars (where the mass ratio ) would lead to a runaway mass transfer and the formation of a common envelope, unless their period is very short (e.g. Paczynski 1976; Tutukov & Yungelson 1988). The common envelope is expected to lead to severe orbital shrinkage, inspired by the observation that many close binaries with white dwarf (WD) components have orbits that must have shrunk a lot (Paczynski 1976). However, many observed post-mass transfer binary systems containing low-mass stars, including barium stars and blue stragglers in clusters, have orbital periods that are relatively wide (and often eccentric) that do not fit with this picture (see Jorissen 2019, and references therein).
@@
Strong orbital shrinkage in the first phase of mass transfer also leads to problems with double WDs that would go through a common-envelope phase twice. The second common envelope would then necessarily happen in a tighter binary than the first and the strict core mass – radius relation of giants with degenerate cores would then produce double WDs where the younger (second) WD was significantly lower in mass than the older (first formed) WD (e.g. Han et al. 1995; Iben et al. 1997; Han 1998), contrary to many observations (Moran et al. 1999; Nelemans et al. 2000).
@@
Two solutions have been put forward for this problem. The first one suggests that the first common envelope is qualitatively different from the second, due to the larger mass ratio. This implies larger angular momentum in the binary system and the phase between the onset of Roche-lobe overflow and the actual engulfment would take longer (if the engulfment happens at all), leading to wider orbits after the interaction. A formalism was proposed to calculate the outcome based on angular momentum, assuming the mass lost from the binary takes away angular momentum per unit mass that is a constant times the average angular momentum per unit mass. This ”” formalism, was found to work well to explain the observations111This is somewhat surprising, since (as people have brought forward as criticism on this approach) the strong sensitivity of the orbital change to could artificially lead to a narrow range of acceptable -values from the observations. However, then using in forward modelling would then lead to very wrong result, contrary to our findings (e.g. Nelemans et al. 2001; Toonen et al. 2012)(Nelemans et al. 2000), for values of .
@@
A second suggestion is that the mass transfer is in fact stable (e.g. Webbink 2008; Woods et al. 2012; Li et al. 2023; Zhang et al. 2024). Webbink (2008) postulates the formation of all double WDs from the small initial range of binaries that interact before having a deep convective envelope. That assumption cannot explain the relatively large number of systems. On the other hand, Woods et al. (2012) find models where mass transfer is stable, despite the convective envelope of the donor. More recently, a number of studies also find that the reaction of the donor to mass loss is much more favourable for stability than thought before (e.g. Ge et al. 2020; Temmink et al. 2023). In addition, some objects are found that fall near the period – WD mass relation expected after stable mass transfer (Parsons et al. 2023). Zhang et al. (2024) study the formation of double helium WDs through stable mass transfer plus common envelope evolution and use that to constrain the common-envelope phase. Li et al. (2023) have investigated the effect of the assumption that the first phase of mass transfer was stable on the population of double WDs and find that it reduces the number of double WD. Double WDs have wider relevance e.g. for type Ia supernovae (e.g. Maoz et al. 2014; Liu et al. 2023; Ruiter & Seitenzahl 2024) and as gravitational wave sources for the LISA mission (e.g. Nelemans et al. 2001; Ruiter et al. 2010; Liu et al. 2010; Yu & Jeffery 2010; Korol et al. 2017; Lamberts et al. 2019; Amaro-Seoane et al. 2023; Tang et al. 2024), so it is important to understand their formation.
@@
In this paper, we investigate if we can observationally test the hypothesis that the first phase of mass transfer generally was stable. Many of the binaries that we observe directly after this phase are difficult to interpret due to very strong selection effects, as they are found through their peculiar abundances or in clusters and the non-evolved star easily outshines the WD. We therefore use observations of systems that have evolved further, via a second mass-transfer phase to become double WD or subdwarf B (sdB) + WD systems. The first phase of mass transfer produces a WD that we observe as the older WD, while the second mass-transfer phase produces the observed younger WD or the sdB. As shown in previous work (e.g. Nelemans et al. 2000; De Marco et al. 2011), the period before the second mass-transfer phase, and thus after the first one can be reconstructed from the observations (also providing information about the common envelope evolution in the second phase of mass transfer, e.g. De Marco et al. 2011; Scherbak & Fuller 2023).
@@ 2 Method
The formation of a close pair of WDs generally proceeds as follows: the most massive star (primary, indicated with subscript 1 below) evolves to become a red giant and at some point in its evolution fills its Roche lobe and starts mass transfer. After the mass transfer, the remaining core of the most massive star becomes a WD (WD1). The initially less massive star (secondary, indicated as star 2 below) evolves later and again starts mass transfer when it fills its Roche lobe, typically on the giant branch and forms a second WD (WD2). Both WD might first be sdB stars for a short period, if their mass is close to 0.47.
For low-mass stars, , the evolution after the main sequence quickly reaches the red giant branch, where these stars develop degenerate helium cores. Due to the tight core mass – luminosity relation of such giants and the fact that they have deep convective envelopes and thus reside on the Hayashi track, there also is a tight core mass – radius relation for such giants. There are a number of different studies that determine the core mass – radius relation (or alternatively the period – core mass relation for giants filling their Roche lobe). Based on these studies, we construct two curves we will use in this paper (see Appendix A for more details), one for approximately Solar metallicity (), one for intermediate metallicity ():
with for Solar metallicity and for intermediate metallicity. These bracket the bulk of the star formation in the Milky Way (see Appendix A). At very low core mass and short period, the core mass - period relation may not always hold (Chen et al. 2017, 2022) and in general this relation is somewhat model dependent (see Appendix A)
Because the core mass – radius relation is very well-defined for giants with degenerate helium cores, we restrict ourselves to systems for which Roche-lobe overflow in this first mass-transfer phase leads to the formation of a helium WD. These have masses in the range between the minimum core mass at the base of the red giant branch () and the maximum core mass, at which the core undergoes the helium flash (). We investigate different scenarios for how this first phase of mass transfer proceeds.
A second phase of mass transfer leads to the formation of a double WD system. Because of the large mass ratio at the onset of this second mass transfer (donor with and accretor with ) as well as the short periods of the current orbits, we assume that the second phase of mass transfer must have been a common-envelope phase with significant shrinkage of the orbit. We also assume that this happened on such a short timescale that the core mass of the giant at the start of the mass transfer is equal to the mass of the second formed WD ().
@@ 2.1 Reconstructing the evolution using the observed masses
We can reconstruct the evolution of the binary based only on the masses by using the core mass – radius relation, together with the fact that the stars that follow this type of evolution have masses in a relatively small range from about 1 to 2. Below 1, the stars do not evolve in the age of the Universe, above 2, the stars do not form degenerate cores on the red giant branch. A very similar approach is taken by Zhang et al. (2024).
@@
Let us denote the initial masses of the two stars as and and the initial orbital period as and initial separation . We use as subscript for the properties after the first mass transfer, but before the second mass transfer. The masses of the WDs formed from the primary and secondary are indicated with and . The latter is observed as the younger, brighter object and thus in observational work is typically referred to as . We neglect initial eccentricity as the orbit is expected to circularize before mass transfer starts. The primary fills its Roche lobe at the moment that
@@
with and the size of the Roche lobe in units of the separation, as given e.g. in Eggleton (1983). In the case of stable mass transfer, the evolution proceeds until the envelope is lost, while the star, at least at the end of the evolution, is in thermal equilibrium (e.g. Rappaport et al. 1995). That means the star still obeys the core mass – radius relation, and thus the period and mass of the remaining WD () after stable mass transfer can be obtained from the condition that the evolution ends with a Roche lobe that is equal to the radius of the giant with core mass equal to . This leads to the well-known WD – period relation that is observed in millisecond radio pulsars (e.g. Rappaport et al. 1995; Tauris & Savonije 1999). Because of the very similar dependence of and on the mass of the companion, this relation is almost independent of and only depends on and metallicity. It can be approximated very well by (Lin et al. 2011)
@@
where the scale factor is 1.1 for Solar metallicity and 0.8 for intermediate metallicity to match our radii (eq. (1)). From the observed mass of the first formed WD () we can thus reconstruct the period after the first phase of mass transfer if that was stable. A similar approach is used by Zhang et al. (2021) to study sdB + helium WD binaries and Zhang et al. (2024) for double WD. Zhang et al. (2024) compare their semi-analytic model (using a slightly different core mass – radius relatio then ours) with detailed stellar evolution calculations and find small a deviation at the lower end of the period/core mass range. In Sect. 4.2 we show that this small difference does not affect our results.
@@
Next, we can use the core mass – radius relation again, but now using to determine the separation at the start of the second phase of mass transfer, because we assume is equal to the core mass at the onset of mass transfer, i.e. when the radius of the secondary was the size of its Roche lobe. This is possible because we assume that the mass transfer is fast (see Nelemans et al. 2000). It leads to:
@@
from which we can obtain via Kepler’s law. In this case, the reconstructed period does depend on , metallicity and the mass of the secondary , because it is the Roche lobe of that star that we calculate. For each system, we simply calculate for the limiting masses of 1 and 2 from the separation given above. We can now check if the reconstructed periods agree with those from stable mass transfer, i.e. for both metallicities. If so, the first phase of mass transfer could have been stable.
@@ 2.2 Forward modelling of the expected masses
As second step, we calculate where in the plane we expect to find close double WDs for three different assumptions for the first phase of mass transfer: common envelope, stable mass transfer, or angular momentum balance using .
We start with a large number (30000) of uniformly randomly selected values for the initial masses, and plus the mass of first formed WD, to sample the possible progenitors, without taking into account the relative likelihood of the different progenitors. We split the sample in two for the metallicity: 2/3 of the points have Solar metallicity and 1/3 have intermediate metallicity (see Appendix A). For each of the three different assumptions we calculate the period after the first mass transfer (). For the case of stable mass transfer we use the equations above to calculate , while for the common envelope and angular momentum balance we calculate by assuming the mass transfer was rapid, so it started when the primary already had a core mass equal to to calculate the initial period. We then use the standard equations to calculate (e.g. Nelemans et al. 2004)
where is the envelope mass fraction, is the common envelope efficiency times a parameter describing the structure of the giant and is the fraction of specific angular momentum lost. From these, the periods after the mass transfer, and are calculated.
The mass of the second WD is now given by the core mass of the giant of mass (we assume no mass is accreted, but that has little influence) that fills its Roche lobe in a binary with period and companion mass .
@@ 2.3 Reconstructing values
Finally, we can also repeat the procedure of Nelemans et al. (2000, 2004), where the observed masses are used to reconstruct both the first and second phase of mass transfer, assuming the first was guided by the angular momentum balance, and the second was a common envelope. For each observed binary the possible values of that would give the right final masses can be calculated for a sample of progenitor masses, that we take uniform between 1 and 2 . We also sample from the uncertainties in the observed masses of the two WDs. Each set of parameters gives rise to a different value of . The same procedure can be used to calculate the possible values of in case the first phase of mass transfer was a common envelope as well.

### v2

@@ Abstract
Context. The masses of the white dwarfs in a binary carry information about previous mass-transfer phases. The core mass – radius relation of low-mass giants gives the size of the progenitor of a helium white dwarf at the moment it last filled its Roche lobe. Previously, we used this information for a few observed systems to propose a new mass-transfer type based on an angular momentum balance.
Aims. Our aim is to investigate if stable mass transfer instead of the angular-momentum prescription is consistent with the observed double-helium white-dwarf masses.
Methods. We reconstructed the progenitor evolution of observed double-helium white dwarfs using the core mass – radius relation and evaluated if the periods at the start of the second phases of mass transfer are consistent with the outcome of stable mass transfer. More generally, we calculated the mass distribution of double-helium white dwarfs for three different progenitor scenarios: double common envelope (with parameter ), angular-momentum prescription (with parameter ), and stable mass transfer.
Results. We find that the observed systems are generally not consistent with stable mass transfer. Stable mass transfer leads to a tight correlation between the two white dwarf masses in a binary that is not consistent with the observed mass distribution. Double-common-envelope evolution is a particularly poor fit to the observations. The angular-momentum prescription can populate the observed mass distribution, but not perfectly.
Conclusions. We conclude that the first phase of mass transfer initiated on the red giant branch in low-mass systems does not generally proceed as stable mass transfer nor as a common envelope, and thus it is poorly understood. This may be related to the fact that for many observed binaries that have finished the first phase of mass transfer the orbit is eccentric, which is an unexpected outcome of mass transfer.
@@
Most stars in the universe have low masses, and all of these low-mass stars (0.8-2) follow a similar evolution (e.g. Refsdal and Weigert 1971; Paczyński 1971): after the main sequence the helium core contracts, becomes degenerate, and grows in mass via hydrogen-shell burning. At the same time, the envelope of the star expands, and the star ascends the giant branch until the core experiences the helium flash around a mass of 0.47. The expansion of these stars happens for a significant part when they have deep convective envelopes.
@@
For a long time, it was believed that the first mass-transfer phase for binary systems with such stars (where the mass ratio ) would lead to a runaway mass transfer and the formation of a common envelope, unless their period is very short (e.g. Paczynski 1976; Tutukov and Yungelson 1988). The common envelope is expected to lead to severe orbital shrinkage, inspired by the observation that many close binaries with white-dwarf (WD) components have orbits that must have shrunk a lot (Paczynski 1976). However, many observed post-mass-transfer binary systems containing low-mass stars, including barium stars and blue stragglers in clusters, have orbital periods that are relatively wide (and often eccentric) that do not fit with this picture (see Jorissen 2019, and references therein).
@@
Strong orbital shrinkage in the first phase of mass transfer also leads to problems with double WDs that would go through a common-envelope phase twice. The second common envelope would then necessarily happen in a tighter binary than the first, and the strict core mass – radius relation of giants with degenerate cores would then produce double WDs where the younger (second) WD is significantly lower in mass than the older (first formed) one (e.g. Han et al. 1995; Iben et al. 1997; Han 1998), contrary to many observed systems (Moran et al. 1999; Nelemans et al. 2000).
@@
Two solutions to this problem have been put forward. The first one suggests that the first common envelope is qualitatively different from the second, due to the larger mass ratio. This implies a larger angular momentum in the binary system and the phase between the onset of Roche-lobe overflow and the actual engulfment would take longer (if the engulfment happens at all), leading to wider orbits after the interaction. A formalism was proposed to calculate the outcome based on angular momentum, assuming that the mass lost from the binary takes away angular momentum per unit mass; that is, a constant times the average angular momentum per unit mass. This ’’ formalism was found to explain the observations well111This is somewhat surprising, since (as people have brought forward as criticism of this approach) the strong sensitivity of the orbital change to could artificially lead to a narrow range of acceptable -values from the observations. However, then using in forward modelling would lead to erroneous results, contrary to our findings (e.g. Nelemans et al. 2001; Toonen et al. 2012).(Nelemans et al. 2000) for values of .
@@
A second suggestion is that the mass transfer is in fact stable (e.g. Webbink 2008; Woods et al. 2012; Li et al. 2023; Zhang et al. 2024). Webbink (2008) postulated the formation of all double WDs from the small initial range of binaries that interact before having a deep convective envelope. That assumption cannot explain the relatively large number of systems. Woods et al. (2012) find models where mass transfer is stable, despite the convective envelope of the donor. More recently, a number of studies also find that the reaction of the donor to mass loss is much more favourable for stability than was previously thought (e.g. Ge et al. 2020; Temmink et al. 2023). In addition, some objects are found that fall near the period – WD mass relation expected after stable mass transfer (Parsons et al. 2023). Zhang et al. (2024) studied the formation of double-helium WDs through stable mass transfer plus common-envelope evolution and used that to constrain the common-envelope phase. Li et al. (2023) investigated the effect of the assumption that the first phase of mass transfer was stable on the population of double WDs and find that it reduces the number of double WD. Double WDs are more widely relevant, for example for type Ia supernovae (e.g. Maoz et al. 2014; Liu et al. 2023; Ruiter and Seitenzahl 2024) and as gravitational wave sources for the LISA mission (e.g. Nelemans et al. 2001; Ruiter et al. 2010; Liu et al. 2010; Yu and Jeffery 2010; Korol et al. 2017; Lamberts et al. 2019; Amaro-Seoane et al. 2023; Tang et al. 2024), so it is important to understand their formation.
@@
In this work, we investigated the possibility of observationally testing the hypothesis that the first phase of mass transfer was generally stable. Many of the binaries that we observe directly after this phase are difficult to interpret due to very strong selection effects, as they are found through their peculiar abundances or in clusters, and the non-evolved star easily outshines the WD. We therefore used observations of systems that have evolved further, via a second mass-transfer phase to become double WD or subdwarf B (sdB) + WD systems. The first phase of mass transfer produces a WD that we observed as the older WD, while the second mass-transfer phase produces the observed younger WD or the sdB. As shown in previous work (e.g. Nelemans et al. 2000; De Marco et al. 2011), the period before the second mass-transfer phase, and thus after the first one, can be reconstructed from the observations (also giving information about the common envelope in the second phase of mass transfer, e.g. De Marco et al. 2011; Scherbak and Fuller 2023).
@@ 2 Method
The formation of a close pair of WDs generally proceeds as follows: the most massive star (primary, indicated with subscript 1 below) evolves to become a red giant and at some point in its evolution fills its Roche lobe and begins mass transfer. After the mass transfer, the remaining core of the most massive star becomes a WD (WD1). The initially less massive star (secondary, indicated as star 2 below) evolves later and again starts mass transfer when it fills its Roche lobe; this typically happens on the giant branch, and then a second WD (WD2) is formed. Both WDs might first be sdB stars for a short period, if their mass is close to 0.47.
For low-mass stars, , the evolution after the main sequence quickly reaches the red giant branch, where these stars develop degenerate helium cores. Due to the tight core mass – luminosity relation of such giants and the fact that they have deep convective envelopes and thus reside on the Hayashi track, there also is a tight core mass – radius relation for such giants. There are a number of different studies that determine the core mass – radius relation (or, alternatively, the period – core mass relation for giants filling their Roche lobe). Based on these studies, we constructed the two curves we use in this paper (see Appendix A for more details): one for approximately solar metallicity () and one for intermediate metallicity ();
with for solar metallicity and for intermediate metallicity. These bracket the bulk of the star formation in the Milky Way (see Appendix A). At a very low core mass and short period, the core mass – period relation may not always hold (Chen et al. 2017, 2022), and in general this relation is somewhat model dependent (see Appendix A).
Because the core mass – radius relation is very well defined for giants with degenerate helium cores, we restricted ourselves to systems for which Roche-lobe overflow in this first mass-transfer phase leads to the formation of a helium WD. These have masses in the range between the minimum core mass at the base of the red giant branch () and the maximum core mass, at which point the core undergoes the helium flash (). We investigated different scenarios of how this first phase of mass transfer may proceed.
A second phase of mass transfer leads to the formation of a double-WD system. Because of the large mass ratio at the onset of this second mass transfer (donor with and accretor with ) as well as the short periods of the current orbits, we assume that the second phase of mass transfer must have been a common-envelope phase with significant shrinkage of the orbit. We also assume that this happened on such a short timescale that the core mass of the giant at the start of the mass transfer is equal to the mass of the second formed WD ().
@@ 2.1 Reconstructing the evolution using the observed masses
We can reconstruct the evolution of the binary based only on the masses by using the core mass – radius relation, together with the fact that the stars that follow this type of evolution have masses in a relatively small range from about 1 to 2. Below 1, the stars do not evolve in the age of the Universe, above 2, the stars do not form degenerate cores on the red giant branch. A very similar approach was taken by Zhang et al. (2024).
@@
Let us denote the initial masses of the two stars as and and the initial orbital period as and initial separation . We use as subscript for the properties after the first mass transfer, but before the second mass transfer. The masses of the WDs formed from the primary and secondary are indicated with and . The latter is observed as the younger, brighter object and thus in observational work is typically referred to as . We neglected initial eccentricity, as the orbit is expected to circularise before mass transfer starts. The primary fills its Roche lobe at the moment that
@@
with and being the size of the Roche lobe in units of the separation, as given, for example, in Eggleton (1983). In the case of stable mass transfer, the evolution proceeds until the envelope is lost, while the star, at least at the end of the evolution, is in thermal equilibrium (e.g. Rappaport et al. 1995). That means the star still obeys the core mass – radius relation, and thus the period and mass of the remaining WD () after stable mass transfer can be obtained from the condition that the evolution ends with a Roche lobe that is equal to the radius of the giant with core mass equal to . This leads to the well-known WD – period relation that is observed in millisecond radio pulsars (e.g. Rappaport et al. 1995; Tauris and Savonije 1999). Because of the very similar dependence of and on the mass of the companion, this relation is almost independent of and only depends on and metallicity. It can be approximated very well by (Lin et al. 2011)
@@
where the scale factor is 1.1 for solar metallicity and 0.8 for intermediate metallicity to match our radii (Eq. (1)). From the observed mass of the first formed WD (), we can thus reconstruct the period after the first phase of mass transfer if that was stable. A similar approach is used by Zhang et al. (2021) to study sdB + helium WD binaries and Zhang et al. (2024) for double WDs. Zhang et al. (2024) compared their semi-analytic model (using a slightly different core mass – radius relation than ours) with detailed stellar-evolution calculations and found a small deviation at the lower end of the period or core mass range. In Sect. 4.2, we show that this small difference does not affect our results.
@@
Next, we can use the core mass – radius relation again, but now using to determine the separation at the start of the second phase of mass transfer, because we assume is equal to the core mass at the onset of mass transfer; that is, when the radius of the secondary was the size of its Roche lobe. This is possible because we assumed that the mass transfer is fast (see Nelemans et al. 2000). It leads to
@@
from which we can obtain via Kepler’s law. In this case, the reconstructed period does depend on , metallicity, and the mass of the secondary , because it is the Roche lobe of that star that we calculated. For each system, we simply calculated for the limiting masses of 1 and 2 from the separation given above. We can now check if the reconstructed periods agree with those from stable mass transfer (i.e. for both metallicities). If so, the first phase of mass transfer could have been stable.
@@ 2.2 Forward modelling of the expected masses
In the second step, we calculated where in the plane we expect to find close double WDs for three different assumptions for the first phase of mass transfer: common envelope, stable mass transfer, or angular momentum balance using .
We start with a large number (30000) of uniformly and randomly selected values for the initial masses, and plus the mass of the first formed WD, to sample the possible progenitors, without taking into account the relative likelihood of the different progenitors. We split the sample in two for the metallicity: 2/3 of the points have solar metallicity and 1/3 have intermediate metallicity (see Appendix A). For each of the three different assumptions, we calculated the period after the first mass transfer (). For the case of stable mass transfer we used the equations above to calculate , while for the common envelope and angular momentum balance we calculated by assuming the mass transfer was rapid, so it started when the primary already had a core mass equal to to calculate the initial period. We then used the standard equations to calculate (e.g. Nelemans and Tout 2005):
where is the envelope mass fraction, is the common-envelope efficiency times a parameter describing the structure of the giant, and is the fraction of specific angular momentum lost. From these, the periods after the mass transfer, and are calculated. The mass of the second WD is now given by the core mass of the giant of mass (we assume no mass is accreted, but that has little influence) that fills its Roche lobe in a binary with period and companion mass .
@@ 2.3 Reconstructing values
Finally, we also repeated the procedure of Nelemans et al. (2000); Nelemans and Tout (2005), where the observed masses are used to reconstruct both the first and second phases of mass transfer, assuming the first was guided by the angular momentum balance and the second was a common envelope. For each observed binary, the possible values of that would give the right final masses can be calculated for a sample of progenitor masses, which we took as uniform between 1 and 2 . We also sampled from the uncertainties in the observed masses of the two WDs. Each set of parameters gives rise to a different value of . The same procedure can be used to calculate the possible values of in case the first phase of mass transfer was a common envelope as well.